Circular No. 7870

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cb IN MCG +8-24-34 T. Boles, Coddenham, Suffolk, reports his discovery of an apparent

supernova (mag 17.0) on a 30-s unfiltered CCD image taken on Apr. 7.145 UT (limiting mag 18.5) with a 0.35-m Schmidt-Cassegrain reflector, in the course of the U.K. Nova/Supernova patrol. The new object is confirmed at a similar magnitude in a subsequent image taken by R. Arbour, South Wonston, Hants., on Apr. 7.851 on a 45-s CCD image using a 0.30-m f/6.3 Schmidt-Cassegrain reflector. It is also confirmed by M. Armstrong, Rolvenden, Kent, on Apr. 7.862 (0.35-m Schmidt-Cassegrain + CCD). The object is located at α = 13^{h}04^{m}23^{s}.97, δ = +47^{o}35′53′′.2, which is 9′′.5 east and 6′′.4 north of MCG +8-24-34. Armstrong gives end figures 23^{s}.92, 53′′.1. The new object is not present on Palomar Sky Survey images taken on 1991.280 (red, limiting mag 20.5) and 1995.236 (blue, limiting mag 22.0).

R. Chornock, W. D. Li and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm) obtained on Apr. 8 with the Shane 3-m reflector at Lick Observatory reveals that the object is indeed a supernova, of type IIn. The spectrum exhibits weak, narrow emission lines of the hydrogen Balmer series superposed on a smooth, almost featureless blue continuum.

XTE J1751-305 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and J. H. Swank, GSFC, report that the orbital period of the recently discovered pulsar XTE J1751-305 (IAUC 7867) is 2545.3(1) s. RXTE PCA observations taken intermittently during Apr. 4.64–7.54 UT show doppler modulation of the pulsed signal, with a projected semimajor axis of 10.06(6) light-ms. The barycentric mean spin frequency of the pulsar is 435.31812(4) Hz, and the epoch of 90^{o} mean longitude is JD 2452369.1389(1). The mass function of the pulsar is 1.26 × 10^{−6}M_{⊙}, which gives a minimum mass for the companion of 0.014M_{⊙} (∼ 15 Jupiter masses). No x-ray eclipses or absorption dips were observed, indicating that the source is not viewed directly edge-on. A power-law fit to the x-ray spectrum has a photon index of 1.9. The neutral-hydrogen-absorption column density for a range of spectral models is 1 to 3×10^{22} cm^{−2}, indicating that the optical counterpart will be heavily absorbed. The x-ray flux is variable at the 1-to-2-percent level (r.m.s.), primarily on the timescale of several orbital periods. The average x-ray flux has declined from 56 to 43 mCrab, corresponding to an exponential-decay constant of 10 days, thus justifying prompt follow-up observations.

2002 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden